# Spectral properties from an efficient analytical representation of the $GW$ self-energy within a multipole approximation

### v1

@@ II.3 MPA for the self-energy
The MPA-$W$ representation of Eq. (10) shows that $\Sigma_{\text{c}}$ can be written as a sum of poles. However, while $W_{\text{c}}^{\text{MPA}}$ is represented by a small number of poles ($n_{W}$), the evaluation of Eq. (10) for each frequency point still requires a large number of matrix multiplications due to the dependence of $P_{m}^{\text{KS}}$, $\Omega_{p}$ and $R_{p}$ on the ${\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}$ indexes. To solve the linearized QP equation in Eq. (6), $\Sigma$ only needs to be evaluated at two frequencies, or one if the analytic expression for the renormalization factor in Eq. (11) is used. However, to obtain spectral properties beyond the QP pole and the renormalization factor, it must be evaluated for a wide frequency range.
To avoid direct evaluation of the self-energy projection for each KS state on a dense frequency grid, $\Sigma_{n{\mathbf{k}}}$ can be modeled as a simple multipole-Padé approximant with a small number of $n_{\Sigma}$ poles that do not depend explicitly on ${\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}$, but are consistent with Eq. (10):
@@
The corresponding renormalization factor is given by:
$(Z^{\text{MPA-}\Sigma}_{n{\mathbf{k}}})^{-1}=-\sum_{p=1}^{n_{\Sigma}}\frac{S_{n{\mathbf{k}}p}}{\xi_{n{\mathbf{k}}p}^{2}}.$ (13)
Therefore, analogous to what is done in MPA-$W$, $\Sigma$ can be computed for a small number of frequency points, that are in turn interpolated to obtain the poles, $\xi_{p}$, and residues, $S_{p}$, of the $\Sigma^{\text{MPA-}\Sigma}$ model. The $\Sigma$ values can be computed, as done here, from Eq. (10).
@@
Figure 1: Example of an MPA asymmetric sampling in the complex frequency plane with two branches close and far from the real axis, with imaginary part $\varpi=\pm 1$ eV (orange circles) and $\varpi=\pm 20$ eV (blue squares), each having 6 points in the positive side and 8 in the negative one distributed according to the linear semi-homogeneous partition of Refs. [39, 40].
@@
As for MPA-$W$, an adequate frequency sampling of $\Sigma$ in the complex plane is essential for obtaining an effective MPA-$\Sigma$ representation. We adopted the same type of inhomogeneous samplings parallel to the real axis used for MPA-$W$. Unlike $W$ (Eq. (9)), $\Sigma$ is not symmetric in $\omega$ and therefore consists of single poles rather than Lorentzians, i.e., paired poles at $z=\pm\Omega_{p}$. For this reason, $\Sigma$ requires sampling along both the positive and negative axes, with a denser sampling in the region with the maximum variability. This corresponds to negative frequencies for the valence states, and positive for the conduction. The sampling is chosen so that it complies with time ordering, having a small positive (negative) imaginary part for energies larger (smaller) than the KS energies, typically of $0.1$ eV. $\Sigma$ has a smoother structure than $W$, it is therefore sufficient to sample it along a single line parallel to the real frequency axis.
The sampling is illustrated in Fig. 1. The parallel sampling is done along the orange line, while the double parallel would use both the orange and the blue points. Despite the need for sampling $\Sigma$ along the negative part of the real axis, sampling it along a single line allows us to use the same number of sampling points as in the double parallel sampling of $W$, typically about 20, for both MPA-$W$ and MPA-$\Sigma$. While the frequency sampling for MPA-$W$ must ensure that the main peaks of $W$ are well reproduced, such as the plasmon peak, setting the MPA-$\Sigma$ sampling can be more challenging as it requires a particularly accurate interpolation around $z=0$ for obtaining accurate QP energies.
@@ II.4 MPA for the Green’s function
One could construct a multipole-Padé representation for the Green’s function, MPA-$G$, using the same type of interpolation used for MPA-$W$ and MPA-$\Sigma$, with the numerical data of $\{z_{j},G(z_{j})\}$. However, with MPA-$\Sigma$ in place, an MPA-$G$ representation can more conveniently be obtained from the Dyson equation in Eq. (3):
$G^{\text{MPA-}\Sigma}_{n{\mathbf{k}}}(z)=\sum_{p=1}^{n_{\Sigma}+1}\frac{P_{n{\mathbf{k}}p}}{z-\epsilon_{n{\mathbf{k}}p}},$ (14)
where the poles $\epsilon_{p}$ and the residues $P_{p}$ are computed as described in the Appendix B (see also Refs. [29, 69, 57, 58]). Notice that $G^{\text{MPA-}\Sigma}$ has one pole more than $\Sigma^{\text{MPA-}\Sigma}$, corresponding to the QP pole. As mentioned in Sec. II.1 and as will be illustrated in Sec. III.3, the remaining poles correspond to satellites that emerge from the poles of $\Sigma^{\text{MPA-}\Sigma}$.
Constructed in this fashion, the poles of $G^{\text{MPA-}\Sigma}$ are not automatically guaranteed to respect the time ordering, although this can be imposed in a second step.
From Eq. (31), it follows that the sum rule of Eq. (8) is obeyed and simplifies to
$\sum_{p}P_{n{\mathbf{k}}p}=1.$ (15)
The renormalization factor defined previously in Eq. (7), whose analytic expressions are given in Eqs. (11) and  (13), is determined from the local derivative of the self-energy. Using Eq. (15) we can now define a more general form, in which $Z$ is given by the residue of the pole with the largest spectral weight, i.e. the residue of the QP pole:
$Z^{\text{MPA-}G}_{n{\mathbf{k}}}=P_{n{\mathbf{k}}q}:\forall p\neq q,|\mathrm{Re}~{}P_{n{\mathbf{k}}p}|<|\mathrm{Re}~{}P_{n{\mathbf{k}}q}|,$ (16)
where $q$ is the index of the QP pole. Notice that even if the definition in Eq. (7) can be evaluated with any desired numerical precision, the expression is only an approximation of the spectral weight of the QP pole and it does not provide a way to compute the spectral weight of satellites. On the other hand, the MPA-$G$ representation naturally provides the spectral weight of all the poles in exact compliance with the sum rule for the number of particles/holes. Therefore, the numerical accuracy of the QP spectral weight of Eq. (16) depends only on the quality of the MPA-$\Sigma$ interpolation. In Sec. III.3 we discuss the adequacy of each of the two $Z$ definitions in the description of the spectral weight of the QP pole.
@@ II.5 Computational details
DFT calculations were performed using the plane-wave Quantum Espresso package [70, 71] with the Perdew-Burke-Ernzerhof (PBE) variant of GGA [72]. We adopted the norm-conserving optimized Vanderbilt pseudopotentials of Ref. [73], with a kinetic energy cutoff for the wave-functions of $70$, $30$, $100$, and $60$ Ry respectively for Na, Si, Cu, and monolayer MoS2, and $85$ Ry for both the NaCl ion-pair and the F2 molecule. The Brillouin zone was sampled with a $16\times 16\times 16$ Monkhorst-Pack grid for Si, Na, and Cu, $18\times 18\times 1$ for the monolayer MoS2 and $\Gamma$-only for NaCl and F2.
The $G_{0}W_{0}$ calculations were performed with yambo [66, 67]. The screened Coulomb potential was computed from $W^{\text{MPA}}$ in Eq. (10), with $n_{W}=8$ for Si and Na, and $n_{W}=12$ for Cu, the same sampling as in Refs. [39] and [40]. Similarly, for MoS2, NaCl and F2 we used a sampling with 8 poles. Since we are considering a MoS2 monolayer, we used the Monte-Carlo based averaging method ($W$-av) for 2D semiconductors, first developed in Ref. [74] and then merged with MPA-$W$ in Ref. [41]. For metals, we used the constant approximation (CA) method [40] to treat the long-wavelength limit of the intraband contributions. Both, $W$-av and CA are methods that can greatly accelerate the ${\mathbf{k}}$-point convergence of $GW$.
@@ III Results
Figure 2: MPA-$\Sigma$ and MPA-$G$ results compared with the corresponding full-frequency (FF) $\Sigma$ and interacting $G$, for semiconducting and metallic materials. The plots show the spectra of a selected valence (yellow) and a conduction (purple) state of Si (left), Na (middle), and Cu (right), with KS energies specified in their respective panels. Panels (a-c) and (d-f) compares the real and imaginary parts of $\Sigma$, and (g-i) the imaginary part of $G$ (g-i), for the respective materials. The solid orange and purple curves indicate the FF results, while the corresponding MPA results are plotted as black dotted (valence) and dashed-dotted (conduction) lines. In the top and center panels, the red (blue) ticks along the horizontal panel borders indicate the distribution of sampling points along the real frequency axis used as input in the MPA-$\Sigma$ interpolation of the given valence (conduction) state.
Figure 3: MPA-$\Sigma$ and MPA-$G$ results compared with the corresponding full-frequency (FF) $\Sigma$ and interacting $G$, for materials with reduced dimensionality. The plots show the spectra of a selected valence (yellow) and a conduction (purple) state of the monolayer MoS2 (left), the NaCl ion-pair (middle) and the F2 (right) molecule, with KS energies specified in their respective panels. As in Fig. 2, MPA results are plotted as black dotted (valence) and dashed-dotted (conduction) lines, while the red (blue) ticks along the horizontal panel borders indicate the distribution of sampling points along the real frequency axis used as input in the MPA-$\Sigma$ interpolation of the given valence (conduction) state.
@@ III.1 Self-energy and Green’s function of prototypical materials
Fig. 2 shows the real and imaginary parts of the $G_{0}W_{0}$ self-energy, i.e. $\mathrm{Re}~{}\Sigma$ (a-c) and $\mathrm{Im}~{}\Sigma$ (d-f), and the imaginary part of the Green’s function, $\mathrm{Im}~{}G$ (g-i), for a selected valence and a conduction state of Si, Na, and Cu. The solid lines give the results computed with a full-frequency approach (FF), serving as a benchmark, and the (dash-) dotted lines, the MPA-$\Sigma$ approach. The FF approach is evaluated on a grid of 2000 frequency points. For MPA-$\Sigma$, 18 frequencies are used for Si and 22 for Na and Cu, corresponding to $n_{\Sigma}=9$ and $n_{\Sigma}=11$, respectively. The sampling frequencies used in the interpolation are indicated with red (valence) and blue (conduction) ticks along the horizontal panel edges.
Figure 4: QP energies and renormalization factors of Si (left: a, d, g), Na (middle: b, e, h) and Cu (right: c, f, i) as a function of their DFT-KS energies. Panels (a-c) show the QP solution obtained from the numerical full-frequency (FF) self-energy using a recursive method. Panels (d-f) show the difference between the FF QP equation and the FF linearized results (red squares) and the analytical MPA-$\Sigma$ solution (blue circles). Panels (g-i) correspond to $\mathrm{Re}~{}Z$ obtained with the FF linearized equation evaluated with Eq. (11) (red squares), the solution of the linearized QP equation within MPA-$\Sigma$ evaluated with Eq. (13) (blue circles), and the spectral weight of the QP pole in the MPA-$G$ representation as defined in Eq. (16) (yellow triangles).
The self-energies of Si and Na have the typical two-pole structure characteristic of systems with a screening potential dominated by a single plasmon pole. As seen in the denominators of Eq. (10), the two $\Sigma$ poles are the result of the plasmon convoluted, respectively, with valence and conduction states. Cu presents a similar picture, but with several plasmon-like poles in $W_{0}$, as can be seen in Ref. [40], that couple with the single-particle poles of $G_{0}$, resulting in a richer structure of $\Sigma^{G_{0}W_{0}}$.
@@
Figure 5: Band structure of Na, Si, monolayer MoS2 and Cu, with a variable width representing the imaginary part of the quasiparticles $\mathrm{Im}~{}[\epsilon_{n{\mathbf{k}}}]$. The colormap is given by $-\text{sign}\left(\mathrm{Re}~{}[\epsilon_{n{\mathbf{k}}}]\right)\mathrm{Re}~{}[Z_{n{\mathbf{k}}}]$, representing the renormalization factor for valence (orange shades) and conduction (purple shades) states.
@@
In the calculations described above, the $\Sigma$ samplings were adapted to each particular system and states. It is however convenient to apply the same general scheme for the MPA-$\Sigma$ sampling of all the states of each system. To this end, we chose the same type of frequency distribution and number of poles that converge MPA-$W$ for each system. In the case of $\Sigma$, the distribution is centered at each DFT-KS energy, with a small asymmetry on the frequency sampling, i.e., 1 or 2 frequency points more on the negative (positive) side for valence (conduction) states. The total number of frequency points is given by $2n_{\Sigma}$, with $n_{\Sigma}=9$ for Si and Na, and $n_{\Sigma}=11$ for Cu.
@@
Fig. 4 shows the QP energies and $\mathrm{Re}~{}Z$ of a set of valence and conduction bands of Si, Na, and Cu, in a wide range of energies and momenta. Panels (a-c) show the $G_{0}W_{0}$ energies found by recursively solving the QP equation in Eq. (5) with the FF representation, used here as a reference. They show the expected $GW$ stretching of the KS bands and, for Si, band gap opening. Panels (d-f) compare the difference between the results of the non-linearized FF and the linearized FF QP equation, $\Delta\varepsilon^{\text{lin-FF}}=\varepsilon^{\text{FF}}-\varepsilon^{\text{lin-FF}}$, (red squares), and the difference between the non-linearized FF and the analytical MPA-$\Sigma$ results, $\Delta\varepsilon^{\text{MPA-}\Sigma}=\varepsilon^{\text{FF}}-\varepsilon^{\text{MPA-}\Sigma}$, (blue circles). For valence states $\Delta\varepsilon^{\text{lin-FF}}$ increases with the distance from the Fermi level, consistent with the general trends for materials (see e.g., Ref. [64]). In contrast, $\Delta\varepsilon^{\text{MPA-}\Sigma}$ is almost zero for most of the valence and conduction states, with the exception of a few quasiparticles with more structure in the self-energy around $\omega=0$. This good agreement between MPA-$\Sigma$ and FF demonstrates the accuracy of the MPA-$\Sigma$ method, even with a simple frequency sampling scheme.
@@
As mentioned in Sec. II.3, solving the linearized QP equation requires $\Sigma$ to be computed for one or two frequency points, whereas MPA-$\Sigma$ requires about 20. While the additional sampling points do increase computational costs, this comes with a significant improvement in accuracy. Critically, MPA-$\Sigma$ also allows for a straightforward evaluation of $\Sigma$ in its full-frequency range and gives access to an analytical representation of $G$.

### v2

@@ II.3 MPA for the self-energy
The MPA-$W$ representation of Eq. (10) shows that $\Sigma_{\text{c}}$, the correlation part of $\Sigma$, can be written as a sum of poles. However, the evaluation of Eq. (10) for each frequency point still requires a large number of matrix multiplications due to the dependence of $\rho_{nm}^{\text{KS}}$, $\Omega_{p}$ and $R_{p}$ on the ${\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}$ indices. To solve the linearized QP equation in Eq. (6), $\Sigma$ only needs to be evaluated at two frequencies, or one if Eq. (12) is used to compute the renormalization factor. However, to obtain spectral properties beyond the QP pole and the renormalization factor, a wide frequency range is needed.
To avoid the direct evaluation of the self-energy projection for each KS state on a dense frequency grid, $\Sigma_{n{\mathbf{k}}}$ can be modeled as a simple multipole-Padé approximant with a small number of $n_{\Sigma}$ poles that do not depend explicitly on ${\mathbf{q}}{\mathbf{G}}{\mathbf{G}}^{\prime}$, but are consistent with Eq. (10):
@@
The corresponding derivative is given by:
$\frac{\partial\Sigma^{\text{MPA-}\Sigma}_{n{\mathbf{k}}}(z)}{\partial z}=-\sum_{p=1}^{n_{\Sigma}}\frac{S_{n{\mathbf{k}}p}}{(z-\xi_{n{\mathbf{k}}p})^{2}}.$ (14)
Therefore, analogously to MPA-$W$, $\Sigma$ can be explicitly computed for a small number of frequency points used to interpolate the $\Sigma^{\text{MPA-}\Sigma}$ model. For each $n{\mathbf{k}}$ (omitted for simplicity), the poles, $\xi_{p}$, and residues, $S_{p}$, are obtained by solving the following system of $2n_{\Sigma}$ equations and variables:
$\Sigma_{\text{c}}^{\text{MPA-}\Sigma}(z_{i})\equiv\sum_{p}^{n_{\Sigma}}\frac{S_{p}}{z_{i}-\xi_{p}}=\Sigma_{c}(z_{i})\text{, }i=1,\ldots,2n_{\Sigma}.$ (15)
@@
Figure 1: Example of an asymmetric MPA-$\Sigma$ sampling in the complex frequency plane with two branches close and far from the real axis, with imaginary part $\varpi=\pm 1$ eV (orange circles) and $\varpi=\pm 20$ eV (blue squares), each having six points in the positive side and eight in the negative one distributed according to the linear semi-homogeneous partition of Refs. [41, 42].
The solution of Eq. (15) is obtained with a procedure similar to the one used in MPA-$W$, in which the correlation part of the MPA-$\Sigma$ model is rewritten in its Padé form, i.e., as a fraction of two polynomials:
$\sum_{p}^{n_{\Sigma}}\frac{S_{p}}{z-\xi_{p}}=\frac{A_{n_{\Sigma}-1}(z)}{B_{n_{\Sigma}}(z)}.$ (16)
The coefficients of the polynomial $B_{n_{\Sigma}}(z)$ can be evaluated from the numerical reference data, $\{z_{i},\Sigma_{c}(z_{i})\}$, using one of the two methods developed in Ref. [41], based on linear algebra and Thiele’s Padé interpolation (see details in Sec. I of Ref. [73]). Moreover, its factorization can be performed using the companion matrix method [41] and is given by
$B_{n_{\Sigma}}(z)=\prod_{p}^{n_{\Sigma}}(z-\xi_{p}).$ (17)
In both methods the sampling points are divided into two sets, used to separate the problem of finding the poles, $\xi_{p}$, from the much simpler problem of finding the residues, $S_{p}$, once the poles are known. Such separation is computationally advantageous since the nonlinear problem of $2n_{\Sigma}$ variables in Eq. (15), is reduced to two problems of size $n_{\Sigma}$, one nonlinear for the poles and the other linear for the residues. Moreover, by first obtaining the poles, it is then possible to apply physical constraints. We impose the time ordering to the complex poles $\xi_{p}=\zeta_{p}+i\varsigma_{p}$, and that they lay in the vicinity of the real frequency axis, as done for MPA-$W$ [41], which results in $\zeta_{p}/\varsigma_{p}<-1$. The residues $S_{p}$ can then be found by solving a simple linear least-squares problem (see details in Sec. I of Ref. [73]).
@@
As for MPA-$W$, an adequate frequency sampling of $\Sigma$ in the complex plane is essential to obtain an effective MPA-$\Sigma$ representation. We adopted the same type of inhomogeneous samplings parallel to the real axis used for MPA-$W$. Unlike $W$ [Eq. (9)], $\Sigma$ is not symmetric in $\omega$ and therefore consists of single poles rather than pairs at $z=\pm\Omega_{p}$. For this reason, $\Sigma$ requires sampling along both the positive and negative axes, with a denser sampling in the region with the maximum variability. This corresponds to negative frequencies for the valence states, and positive for the conduction. A illustrated in Fig. 1, the sampling is chosen so that it complies with time ordering, having a small positive (negative) imaginary part for energies larger (smaller) than the KS energies, typically of $\varpi=\pm 0.1$ eV. The parallel sampling is done along the orange line, while the double parallel would use both the orange and the blue points. Since $\Sigma$ has a smoother structure than $W$, it is sufficient to sample it along a single line parallel to the real frequency axis.
The pole structure of $\Sigma$ is expected to resemble that of $W$, therefore the same sampling distribution can be used for $\Sigma$, as long as it is replicated on the negative side of the imaginary axis, as illustrated in the example of Fig. 1. The use of an even number of points including the origin results in an asymmetric distribution. Despite the need to sample $\Sigma$ along both the positive and negative parts of the real axis, using a single line allows us to use around the same number of sampling points as in the double-parallel sampling of $W$, typically about 20, for both MPA-$W$ and MPA-$\Sigma$. More details on the MPA-$\Sigma$ sampling and the convergence with the number of poles $n_{\Sigma}$ can be found in Sec. II of Ref. [73].
MPA-$\Sigma$ requires a particularly accurate interpolation around $z=0$ for obtaining accurate QP energies, which sometimes requires a more precise sampling. This can be done by benchmarking the sampling for one or a few selected QP states against the corresponding FF calculations. The sampling can then be replicated for the remaining QPs, without extending the FF calculations. As in the case of MPA-$W$, the computational cost of the $\Sigma$ evaluation can be compared in terms of the number of frequency points for which $\Sigma$ is explicitly evaluated. Therefore, given a FF grid spacing of $\Delta\omega=0.1$ eV in a frequency interval of $100$ eV, MPA-$\Sigma$ is typically around 50 times more efficient than the FF $\Sigma$ evaluation.
@@ II.4 MPA for the Green’s function
As in MPA-$W$ and MPA-$\Sigma$, one could construct a multipole-Padé representation of the Green’s function from the interpolation of the numerical data, $\{z_{i},G(z_{i})\}$. However, with MPA-$\Sigma$ in place, it is more convenient to obtain an MPA-$G$ representation from the Dyson equation in Eq. (3) (see also Refs. [31, 74] and the algorithmic-inversion-method in a sum-over-poles (AIM-SOP) representation of Refs. [59, 60, 62]).
Given Eq. (16), the total MPA-$\Sigma$ can be written in its Padé representation as
$\Sigma^{\text{MPA-}\Sigma}(z)=\frac{\Sigma_{x}B_{n_{\Sigma}}(z)+A_{n_{\Sigma}-1}(z)}{B_{n_{\Sigma}}(z)}.$ (18)
Notice that if we apply physical constraints after the factorization in Eq. (17), and fit the residues thereafter, we will need to reconstruct both the $A_{n_{\Sigma}-1}(z)$ and $B_{n_{\Sigma}}(z)$ polynomials from the new poles and residues using Eq. (18), which is straightforward. We can then obtain MPA-$G$ as
$G^{\text{MPA-}\Sigma}(z)$ $\equiv\frac{B_{n_{\Sigma}}(z)}{C_{n_{\Sigma}+1}(z)},$ (19)
$C_{n_{\Sigma}+1}(z)$ $=zB_{n_{\Sigma}}(z)-\Sigma_{x}B_{n_{\Sigma}}(z)-A_{n_{\Sigma}-1}(z).$
The poles of $G^{\text{MPA-}\Sigma}$ will then correspond to the zeros of $C_{n_{\Sigma}+1}(z)$. Analogous to $B_{n_{\Sigma}}(z)$, the $C_{n_{\Sigma}+1}(z)$ polynomial can be factorized, e.g., using the companion matrix method [41]:
$C_{n_{\Sigma}+1}(z)=\prod_{p}^{n_{\Sigma}+1}(z-\epsilon_{p}),$ (20)
while the residues can be computed using the residue theorem. Constructed in this fashion, the poles of $G^{\text{MPA-}\Sigma}$ do not necessarily respect the time ordering, although this can be imposed in a second step.
The resulting multipole-Padé representation of $G$ for each state $n{\mathbf{k}}$ is then given by
$G^{\text{MPA-}\Sigma}_{n{\mathbf{k}}}(z)=\sum_{p=1}^{n_{\Sigma}+1}\frac{Z_{n{\mathbf{k}}p}}{z-\epsilon_{n{\mathbf{k}}p}},$ (21)
where the poles are obtained (see also Ref. [59]) as
$Z_{n{\mathbf{k}}p}=\frac{\prod_{i}^{n_{\Sigma}}(\epsilon_{n{\mathbf{k}}p}-\xi_{n{\mathbf{k}}i})}{\prod_{i\neq p}^{n_{\Sigma}+1}(\epsilon_{n{\mathbf{k}}p}-\epsilon_{n{\mathbf{k}}i})}.$ (22)
Notice that $G^{\text{MPA-}\Sigma}$ has one pole more than $\Sigma^{\text{MPA-}\Sigma}$, corresponding to the QP pole. As mentioned in Sec. II.1 and as will be illustrated in Appendix. A, the remaining poles correspond to satellites that emerge from the poles of $\Sigma^{\text{MPA-}\Sigma}$. All the QP and satellites $\epsilon_{n{\mathbf{k}}p}$ poles are solutions of the QP equation in Eq. (5).
The MPA-$G$ representation overcomes the limitations of the approximation introduced by the linearized QP equation in Eq. (6), for both the positions and the residues of the QP poles. It also preserves important analytical properties of the solutions of non-linear eigenvalue equations with rational self-energy potentials [75]. From Eq. (22), it follows that
$Z_{n{\mathbf{k}}p}=\left[1-\frac{\partial\Sigma_{n{\mathbf{k}}}^{\text{MPA-}\Sigma}(z)}{\partial z}\bigg{|}_{z=\epsilon_{n{\mathbf{k}}p}}\right]^{-1},$ (23)
while the sum rule of Eq. (8) is obeyed and simplifies to
$\sum_{p}Z_{n{\mathbf{k}}p}=1.$ (24)
Therefore, the numerical accuracy of the QP and satellites spectral weights of Eq. (22) depends only on the quality of the MPA-$\Sigma$ interpolation, while always comply exactly with the sum rule for the number of particles and holes.
@@ II.5 Computational details
DFT calculations were performed using the plane-wave Quantum Espresso package [76, 77] with the Perdew-Burke-Ernzerhof (PBE) variant of the generalized gradient approximation (GGA) [78]. We adopted the norm-conserving optimized Vanderbilt pseudopotentials of Ref. [79], with a kinetic energy cutoff for the wave-functions of $70$, $30$, $100$, and $60$ Ry respectively for Na, Si, Cu, and monolayer MoS2, and $85$ Ry for both the NaCl ion-pair and the F2 molecule. The Brillouin zone was sampled with a $16\times 16\times 16$ Monkhorst-Pack grid for Si, Na, and Cu, $18\times 18\times 1$ for the monolayer MoS2 and $\Gamma$-only for NaCl and F2.
The $G_{0}W_{0}$ calculations were performed with yambo [70, 71]. In all the cases, the screened Coulomb potential was computed within MPA-$W$, using Eq. (9) with $n_{W}=8$ for Si and Na, and $n_{W}=12$ for Cu, the same sampling as in Refs. [41, 42]. Similarly, for MoS2, NaCl and F2 we used a sampling with 8 poles and a linear distribution. The method to evaluate the self-energy $\Sigma$, using full-frequency or the new MPA-$\Sigma$ method, is specified in each case. Since we are considering a MoS2 monolayer, we used the Monte-Carlo based averaging method ($W$-av) for 2D semiconductors, first developed in Ref. [80] and then merged with MPA-$W$ in Ref. [43]. For metals, we used the constant approximation (CA) method [42] to treat the long-wavelength limit of the intraband contributions. Both, $W$-av and CA are methods that can greatly accelerate the ${\mathbf{k}}$-point convergence of $GW$.
Figure 2: MPA-$\Sigma$ and MPA-$G$ results compared with the corresponding full-frequency (FF) $\Sigma$ and interacting $G$, for semiconducting and metallic materials. The plots show the spectra of a selected valence (yellow) and a conduction (purple) state of Si (left), Na (middle), and Cu (right), with KS energies relative to the valence band minimum (VBM), as specified in their respective panels. (a)-(c) and (d)-(f) Compares the real and imaginary parts of $\Sigma$, and (g)-(i) the imaginary part of $G$, for the respective materials. The solid orange and purple curves indicate the FF results, while the corresponding MPA-$\Sigma$ and MPA-$G$ results are plotted as black dotted (valence) and dashed-dotted (conduction) lines. In the top and center panels, the red (blue) ticks along the horizontal panel borders indicate the distribution of sampling points along the real frequency axis used as input in the MPA-$\Sigma$ interpolation of the given valence (conduction) state.
Figure 3: MPA-$\Sigma$ and MPA-$G$ results compared with the corresponding full-frequency (FF) $\Sigma$ and interacting $G$, for materials with reduced dimensionality. The plots show the spectra of a selected valence (yellow) and a conduction (purple) state of the monolayer MoS2 (left), the NaCl ion-pair (middle), and the F2 (right) molecule, with KS energies relative to the valence band minimum (VBM), as specified in their respective panels. As in Fig. 2, the MPA-$\Sigma$ and MPA-$G$ results are plotted as black dotted (valence) and dashed-dotted (conduction) lines, while the red (blue) ticks along the horizontal panel borders indicate the distribution of sampling points along the real frequency axis used as input in the MPA-$\Sigma$ interpolation of the given valence (conduction) state.
@@ III.1 Self-energy and Green’s function of prototypical materials
Figure 2 shows the real and imaginary parts of the $G_{0}W_{0}$ self-energy, i.e. $\mathrm{Re}~{}\Sigma$ (a)-(c) and $\mathrm{Im}~{}\Sigma$ (d)-(f), and the imaginary part of the Green’s function, $\mathrm{Im}~{}G$ (g)-(i), for a selected valence and a conduction state of Si, Na, and Cu. The solid lines give the results computed with a full-frequency evaluation of $\Sigma$ and $G$ (FF), serving as a benchmark, and the (dashed) dotted lines, the MPA-$\Sigma$ approach. The FF approach is evaluated on a homogeneous grid of 2000 frequency points. For MPA-$\Sigma$, 18 frequencies are used for Si and 22 for Na and Cu, corresponding to $n_{\Sigma}=9$ and $11$, respectively. The sampling frequencies used in the interpolation are indicated with red (valence) and blue (conduction) ticks along the horizontal panel edges.
Figure 4: QP energies and renormalization factors of Si [left: (a), (d), (g)], Na [middle: (b), (e), (h)] and Cu [right: (c), (f), (i)] as a function of their KS energies. (a)-(c) Show the QP solution obtained from the numerical full-frequency (FF) self-energy using a recursive method. (d)-(f) Show the difference between the FF QP equation and the FF linearized results (red squares) and the analytical MPA-$\Sigma$ solution (blue circles). (g)-(i) Correspond to the linearized $\mathrm{Re}[Z^{\text{lin}}]$ obtained with the FF MPA-$W$ approach of Eq. (12) (red squares) and the MPA-$\Sigma$ representation of Eq. (14) (blue circles), and the non-linearized $\mathrm{Re}[Z]$ computed as the residue of the QP pole in the MPA-$G$ representation of Eq. (22) (yellow triangles).
The self-energies of Si and Na have the typical two-pole structure characteristic of systems with a screening potential dominated by a single plasmon pole. As seen in the denominators of Eq. (10), the two $\Sigma$ poles are the result of the plasmon convoluted, respectively, with valence and conduction states. Cu presents a similar picture, but with several plasmon-like poles in $W_{0}$, as can be seen in Ref. [42], coupled with the single-particle poles of $G_{0}$, resulting in a richer structure of $\Sigma$.
@@
Figure 5: Band structure of Na, Si, monolayer MoS2, and Cu, with a variable width representing the imaginary part of the quasiparticles $\mathrm{Im}[\epsilon_{n{\mathbf{k}}}]$. The color map is given by $-\text{sign}\left(\mathrm{Re}[\epsilon_{n{\mathbf{k}}}]\right)\mathrm{Re}[Z_{n{\mathbf{k}}}]$, representing the renormalization factor for valence (orange shades) and conduction (purple shades) states.
@@
In order to obtain an accurate MPA-$\Sigma$ representation while preserving the physical meaning of the main poles, in the above calculations, the number of sampling points of $\Sigma$ at each side of the frequency axis was adapted to each particular system and state, using the FF calculations as a reference. It is however convenient to establish a general sampling scheme that does not require a FF reference and can be applied to all the states of each system. For this purpose, we select a frequency distribution and number of poles similar to the one used for MPA-$W$. As mentioned in Sec. II.3, in the case of $\Sigma$, the distribution is centered on each KS energy, with a small asymmetry on the frequency sampling, i.e., 1 or 2 frequency points more on the negative (positive) side for valence (conduction) states. The total number of frequency points is given by $2n_{\Sigma}$, with $n_{\Sigma}=9$ for Si and Na, and $n_{\Sigma}=11$ for Cu.
@@
Figure 4 shows the QP energies and $\mathrm{Re}~{}Z$ of a set of valence and conduction bands of Si, Na, and Cu, in a wide range of energies and momenta. Figures 4(a)-4(c) show the $G_{0}W_{0}$ energies found by recursively solving the QP equation in Eq. (5) with the FF $\Sigma$ representation, used here as a reference. They show the expected $GW$ stretching of the KS bands and, for Si, a band gap opening. Figures 4(d)-4(f) compare the difference between the results of the non-linearized FF and the linearized FF QP equation, $\Delta\varepsilon^{\text{lin-FF}}=\varepsilon^{\text{FF}}-\varepsilon^{\text{lin-FF}}$, (red squares), and the difference between the non-linearized FF and the analytical MPA-$\Sigma$ results, $\Delta\varepsilon^{\text{MPA-}\Sigma}=\varepsilon^{\text{FF}}-\varepsilon^{\text{MPA-}\Sigma}$, (blue circles).
@@
For valence states $\Delta\varepsilon^{\text{lin-FF}}$ increases with the distance from the Fermi level, consistent with the general trends for materials (see, e.g., Ref. [67]). In contrast, $\Delta\varepsilon^{\text{MPA-}\Sigma}$ is almost zero for most of the valence and conduction states, with the exception of a few quasiparticles with more structure around $\Sigma(\omega=\varepsilon^{\text{KS}})$. This good agreement between MPA-$\Sigma$ and FF demonstrates the accuracy of the MPA-$\Sigma$ method, even with a simple frequency sampling scheme. As mentioned in Sec. II.3, solving the linearized QP equation requires $\Sigma$ to be computed for one or two frequency points, whereas MPA-$\Sigma$ requires about 20. While the additional sampling points do increase computational costs, this comes with a significant improvement in accuracy. Critically, MPA-$\Sigma$ also allows for a straightforward evaluation of $\Sigma$ in its full-frequency range and gives access to an analytical representation of $G$.